OpenStudy (anonymous):

f(x)= (37−2x)^1/2 find the domain of the given function

OpenStudy (anonymous):

\(f(x)=\sqrt{37-2x}\) Is that it?

OpenStudy (anonymous):

Like is that the function?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

pls also tell how did you arrive at the answer.

OpenStudy (anonymous):

Ok. So then we know that \(\sqrt{0}=0\). But the "thing" inside the square root cannot be negative. So we have: \[(37-2x) \geq 0\] \[37 \geq 2x\] \[\frac{37}{2} \geq x\] Therefore, the domain is: \[D=\Big[{x \in R} \phantom{.} \Big| \frac{37}{2} \geq x\Big]\]

OpenStudy (anonymous):

hmm... lemme plug this answer in my homework. I have tried this question 5 times, hadn't got it right so far. Hope it works this time.

OpenStudy (anonymous):

I hope so too!

OpenStudy (anonymous):

Nope it says the answer is wrong.

OpenStudy (anonymous):

I had to write it in interval format so I wrote it as [37/2,infinity)

OpenStudy (anonymous):

Oh I see yeah that's another way of writing it: \[D=\left[{x \in R} \phantom{.} \Big| x \in \left(-\infty,\frac{37}{2}\right]\right]\]

OpenStudy (anonymous):

I see i was writing it wrong. thanks dude.

OpenStudy (anonymous):

Anytime man